OpenStudy (loser66):

The remainders when a polynomial P(x) is divided by x-2 and x+1 are respectively 3 and -2. The remainder when P(x) is divided by (x^2-x-2)is K(x). What is the value of K(-4) ?

OpenStudy (loser66):

@zzr0ck3r Please.

OpenStudy (loser66):

@dan815 @ikram002p

OpenStudy (anonymous):

yes please

OpenStudy (ikram002p):

for tomorrow xD

OpenStudy (loser66):

It's ok, friend. It is not my problem, it is from other asker, I just help him to post it here because I don't know how to solve. :(

OpenStudy (ikram002p):

Hint:- (x-2) (x+1)=x^2-x-2 so The remainder when P(x) is divided by (x^2-x-2)is K(x) ={The remainders when P(x) is divided by x-2) *{The remainders when P(x) is divided by x+1) k(X)=3*-2

OpenStudy (loser66):

hey, go to bed, my friend. It is not that. K(x) =ax +b, it is a first degree polynomial.

OpenStudy (ikram002p):

i should do that lol

OpenStudy (loser66):

hehehe.. I am using group theory to solve it or remainder theorem, but it sounds awkward hehehe... not know how to solve it yet.

OpenStudy (loser66):

\(P(x) = a_nx^n+a_{n-1}x^{n-1}+...............+a_0\) what??? I saw @electrokid really???

OpenStudy (anonymous):

yo

OpenStudy (loser66):

hey, is it you?? @electrokid or the net is crazy??

OpenStudy (loser66):

wwwwwwoaaaaaaah, how are you?? @Francesco1907 Sorry friend, I am so excited now, can't think anything.

OpenStudy (anonymous):

ok still waiting for the answer

OpenStudy (anonymous):

isnt this the remainder theorem? If remainder of division by polynomial \(P(x)\) by \((x-a)\) is \(b\), then, \(P(a)=b\).

OpenStudy (loser66):

Yes, it is.

OpenStudy (loser66):

P(2) =3 P(-1) =-2 so??

OpenStudy (anonymous):

P(x) = (x^2 - x- 2)Q(x) + K(x) P(2) = K(2) =3 P(-1) = K(-1) = -2 K(x) = ax + b So, 2x + b = 3 -1x + b = -2 solve these to get K(x)

OpenStudy (anonymous):

so, we follow that \[P(2)=3\] and \[P(-1)=-2\] \[P(x)=(x-2)(x+1)K(x)\] so, \[K(x)=(3)(-2)\] i.e., indepemdemt of x

OpenStudy (loser66):

@electrokid I don't think so. \(P(x) = M(x) (x+2)(x-1) + K(x) \)

OpenStudy (loser66):

@adxpoi I think you are right.

OpenStudy (loser66):

but the equations are 2a+b =3 -a+b = -2

OpenStudy (anonymous):

Yup, @adxpoi got it :)

OpenStudy (loser66):

@Francesco1907 take it, please. You have to do the rest.

OpenStudy (anonymous):

Can I ask 1 more question

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

Yeah sorry my bad, replace x by a in my last post :P

OpenStudy (loser66):

@Francesco1907 post a new one by yourself.

OpenStudy (loser66):

hey, it is 0, I think and it is not hard to see it, right? just take derivative and plug pi/10 in.

OpenStudy (anonymous):

nope we found it its -5 mate

OpenStudy (loser66):

how? show me. Don't forget chain rule

OpenStudy (anonymous):

solve me 57 while I am solving your question

OpenStudy (loser66):

I got -1/5

OpenStudy (anonymous):

thats not the correct answer mate

OpenStudy (amilapsn):

For the earlier question K(4) can be found using this method too.. (Using similar method you can find K(-4))